OpenStudy (anonymous):

integrate 2t +(2/t) from 1 to e

OpenStudy (campbell_st):

\[\int\limits_{1}^{e} 2t + 2/t dt = [t^2 + 2\ln(t)] \] so \[f(t) = t^2 + 2\ln(t)\] evaluate by finding f(e) - f(1)

OpenStudy (anonymous):

could you show me... please

OpenStudy (campbell_st):

well f(e) = e^2 + 2 x ln(e) \[f(e) = e^2 + 2 \times 1\] \[f(x) = e^2 + 2\] f(1) = 1^2 + 2ln(1) f(1) = 1 then f(e) - f(1) = e^2 + 2 - 1 the value of the definite integral is e^2 + 1

OpenStudy (anonymous):

arigato!